# Ion Coulomb crystals: an exotic form of condensed matter

### v1

@@ II.3.1 Ground state configuration
More surprisingly, crystal structures mimicking those of infinitely long systems that are harmonically confined in the radial direction can be realized in a linear rf quadrupole [147] (and as well in Penning traps) when two different ion species are simultaneously confined. See the paragraph on Two-species Coulomb crystals which concludes this subsection.
@@
Slab geometry— The slab geometry consists of a uniform density OCP that is harmonically confined in one direction, $q\phi_{T}(\rho,z)=m\omega_{z}^{2}z^{2}/2$, but infinite in the orthogonal directions. With the assumption that the lattice planes that form in the $T=0$ limit have the same symmetry, this geometry allows for a careful theoretical treatment [91] that illustrates the importance of surface effects with 3D trapped ion crystals. Although bcc has a lower bulk energy per ion than fcc, an fcc lattice has a smaller surface energy contribution than bcc, making it the lower energy lattice for thicknesses with smaller numbers of planes. Dubin showed that a slab thickness corresponding to approximately 60 planes is required for bcc to take over from fcc as the prevailing lattice structure with lower energy.
@@
Experimentally this geometry can be approximated in ion traps by adjusting the radial confinement of the trap to be much weaker than the axial confinement. In a Penning trap this condition can be achieved with low rotation frequencies $\omega_{r}$ that are slightly higher than the magnetron frequency $\omega_{m}$ of the trap (see Eq. (24) and Fig. 7). For a given rotation frequency, this is the regime that gives rise to single plane crystals (see Sec. II.2) for a sufficiently small number of ions. With larger numbers of trapped ions, multi-plane crystals form with an overall pancake (or lenticular) shape. The lattice planes have some curvature, but near the radial center ($\rho\approx 0$) of the trap, the planes are very flat and can be used to compare with the predicted structure of infinite planar systems analyzed by [94].
@@
Figure 32: Axial positions of the planes in the central region ($\rho\approx 0$) of small aspect ratio ion crystals as a function of the central areal number density. The lines show predictions from theory, and the symbols show experimental measurements. An open square (filled circle) indicates the experimental observation of a square or rhombic plane (hexagonal plane) was experimentally observed. Similarly, a thin line (thick line) indicates that theory predicts a square or rhombic plane (hexagonal plane). Lengths have been normalized by $a_{WS2D}=(3e^{2}/4\pi\epsilon_{0}m\omega_{z}^{2})^{1/3}$. From 214, where the interlattice displacement vector $\bf{c}_{2}$ is defined.
@@
Figure 32 summarizes the ion crystal structures observed in the laboratory as a function of the 2D areal number density ($\sigma_{0}a_{ws2D}^{2}$) obtained by integrating over the axial extent of the crystal [214]. The crystal structure was determined from direct images of the ion fluorescence from the central ($\rho\approx 0$) region of Doppler laser-cooled ion crystals with aspect ratios $\alpha\equiv Z_{p}/R_{p}<0.1$ and ion number $N\lesssim 10^{4}$ ions. Starting with low 2D number density, single-plane crystals with a triangular lattice (the triangular lattice of Sec.  II.2.2) was observed. With increasing $\sigma_{0}a_{ws2D}^{2}$, theory predicts a “buckling-transition” to three closely spaced planes that is stable over a very small interval of $\sigma_{0}a_{ws2D}^{2}$ (see Sec. II.2.4). This small region of three planes is followed by two, more widely separated planes. Initially each plane has a square and then rhombic symmetry. As the planes move further apart, they transition into planes with hexagonal symmetry. With increasing $\sigma_{0}a_{ws2D}^{2}$, eventually 3 planes with square symmetry form, and the pattern repeats. Fig. 32 also compares the measured separation of the planes with the theory of [94]. The overall agreement between experimental observations with the $T=0$, infinite extent theory of [94] is good, with some minor discrepancies [214]. One disagreement is that the transition from a single plane to a small region of closely spaced three planes was not yet observed. This transition is particularly interesting (see Sec. II.2.4) and motivates a more careful study at lower temperatures with new experimental tools.
@@
Interestingly, the rhombic phases of the few plane ion crystals discussed in the previous paragraph are predicted to have negative Poisson ratios. A negative Poisson’s ratio means that stretching the crystal in a particular direction, in this case, along the $\hat{z}$ or axial direction of the trap, causes an expansion in a lateral direction. This behavior contradicts common experience for ordinary materials that act as incompressible, such as rubber bands. This prediction was verified through experimental measurement [16].
@@
We close this section on slab geometries by pointing out that multilayer 3D crystals of trapped ions consisting of flat planes can be formed through the addition of anharmonic terms to the trapping potential. This was demonstrated theoretically for Penning traps where the anharmonic terms are additions to the usual harmonic trapping potential (see Eq. 20) and are readily implementable in the lab [138]. The formation and control of ion crystal in traps with significant anharmonic components to the trapping potential is a subject for experimental investigation.
@@
Two-species Coulomb crystals— As indicated previously, two-species Coulomb crystals constitute a very interesting and special type of Coulomb crystal. In comparisons to normal solids formed through electronic (chemical) bindings between the atomic constituencies, leading to the formation of various alloys when more atomic species are involved, in Coulomb crystals the confined ions are mainly interacting through the repulsive Coulomb forces at a distance that inhibits chemical bonding. Consequently, the structural organization of two simultaneously trapped ion species will be one that minimizes the sum of the potential energy of all ions. Since in both linear Paul (rf) traps and Penning traps, the effective confining potential in the plane perpendicular to the static confining direction depends on both the charge and mass of a specific ion species, a two-species Coulomb crystal will generally lead to a separation of the two species with respect to this plane. This type of separation has been observed in both types of traps, where the strongest bound species forms a cylinder-like central structure within an ellipsoidal, globally shaped crystal [147, 228]. Examples of such two species crystals are presented in Fig. 33 and Fig. 34. Only in the specific cases when the charge-to-mass ratio of the two species is the same, will they mix as has been observed experimentally in experiments with $^{26}Mg^{+}$ and $^{24}MgD^{+}$ ions [217], as well as in simulations of two-species ion systems where one species has both double the mass and charge of the other species [202]. In the latter case it was found that the ground state of an infinite system would be a simple cubic (SC) lattice with the unit cell consisting of one species at a corner and the other displaced half way the $\langle 111\rangle$ diagonal (like CsCl).
@@
Figure 33: Sketch of a linear rf trap experiment to study two-species ion Coulomb crystals consisting of Mg+ and Ca+ ions (a). Fluorescence images of two-species consisting of 300 Mg+ ions (red) and 3000 Ca+ ions (blue)(b), and 47 Mg+ ions and 1300 Ca+ ions (c). In reality, both crystals have a spheroidal global shape with a cylindrical symmetry axis coinciding with the central axis of the Mg+ component. From 147.
@@ II.3.2 Metastable and perturbation induced structures
In the previous section we have assumed the effective confining potential to be purely harmonic in all directions. This assumption is, though, only an approximation to real experimental situations, where details of trapping potential can produce deviations from this idealized situation and impact the formation of ion crystal structures.
@@
For instance, in the case of linear rf traps operated with $a_{dc}=0$ we have a pseudo potential that has the same strength in the plane perpendicular to the $z$-axis. However, due to the four electrode configuration of the trap (See Fig. 4), the actual ion motional dynamics including the micromotion must have a four fold symmetry with respect to rotation around the $z$-axis. The effect of this breaking of cylindrical symmetry can already be observed when considering only two ions forced by a strong confining force along the $z$-axis to lie in the $xy$-plane. From MD simulations as well as expansion of the pseudo potential beyond the dominating order [86], one finds that the two ions always align along one of the (1,1) or (1,-1) directions in the $xy$-plane, where the orientation of the micromotion is perpendicular to the line connecting the two ions. Similarly, for weak confinement along the $z$-axis, the crystal plane formed by smaller two-dimensional zig-zag structures will align in the same directions in the $xy$-plane. The effect of non-perfect cylindrical symmetry is much more challenging to predict for three-dimensional crystals, since in this case micromotion will happen along many directions.
@@
A general observation is, though, that three-dimensional crystals (being multi-species or not) do not rotate around the $z$-axis when laser-cooled only along the $z$-axis. In experiments where translational crystal structures have been observed within single species crystals, there has been a tendency to always observe, e.g., fcc or bcc structures at specific orientations with respect to the trap electrodes [227, 228], indicating breaking of cylindrical symmetry. Even more pronounced symmetry breaking effects have been observed in two-species crystals consisting of two isotopes of Ca+ (40Ca+ and 44Ca+) [228] (See Fig. 34). Here, the inner 40Ca+ component, in addition to having a global cylindrical shape, had a very persistent (life times longer than 10 s) internal structure (presumable orthorhombic) with fixed relation to the trap electrodes. These findings were to some extent confirmed by MD simulations of similar two-species systems including the full rf field configuration[228].
@@
Figure 34: Fluorescence images of two-species ion Coulomb crystals containing $\sim$1500 $^{40}Ca^{+}$ ions (red) and $\sim$2000 $^{44}Ca^{+}$ ions (blue) for the same ion densities, but different aspect ratios of the crystals. From  228.
Whereas the structural effects discussed above originates from small perturbations governed by the global trapping fields, more spatially localized perturbations to the ions’ trapping potential can be introduced by the application of optically induced dipole forces. For instance, near resonant standing wave light fields might be used to create a one-dimensional periodical potential with local potential minima separated by $\lambda/2$, into which ions can be forced to be localized [197, 104, 188] (See also Sec. III.5). By having the two-dimensional equipotential planes of such a standing wave field made commensurable with particular lattices planes of bcc, fcc, or hcp Coulomb crystals, crystal structures and orientations might potentially be controlled by such perturbations [145]. MD simulations presented in Ref. [145], showed not only that it seems possible to realize such control experimentally, but also indicated the option of adiabatically changing the crystal structure from bcc to fcc structures, and vice versa. Another light induced perturbation that can be imposed on a crystal is a torque produced by the radiation force from a near resonant light beam (as the ones used for laser cooling) that does not propagate through the center of the crystal. By applying such a laser beam in the plane perpendicular to the $z$-axis of a linear rf trap, it has been demonstrated that crystals can be forced to rotate around the $z$-axis with rotational frequencies varying from zero all the way up to 98% of the trap confining frequency in the xyx-y-plane [88] (See Fig. 35).
@@
A more dramatic kind of perturbation one can induce is when in an rf trap one increases the density of a crystal by appropriate increase of the stability parameter $q$ [147] to the point, where the plasma frequency of the crystals becomes close to the half that of the rf drive frequency. At this instance, the general rf field heating [26] of the crystal will increase very rapidly due to parametric resonance and in many instances lead to a fast de-crystallization. However, under appropriately laser cooling conditions, dynamical stable crystal structures with ions locked into seemingly rotating discs can be formed [176](See Fig. 36). Though a full understanding of these types of crystals is still not available, there seems to be no doubt that the stability of these structures are based on a detailed balance between the laser cooling rate and the parametric heating rate. The latter will change significantly when the crystal density, and hence the plasma frequency, lowers due to the rotation, effectively bringing the crystal out of parametric resonance.
@@
As mentioned in Sec. II.3.1, the number of metastable configurations within a given energy interval increases typically exponentially with the number of ions in a Coulomb crystal. Hence, in experiments where the ions are cooled to finite temperatures, one will expect to observe thermally excited metastable configurations. Fig. 37 shows four images recorded seconds apart of the same cold ion ensemble of approximately 2700 ions stored in an rf trap. In this relatively small ensemble, the ion crystal configuration is quickly changing between different 3D periodic lattices, which are excited metastable configurations. We note that Coulomb crystals as small as $\sim$1000 ions have been observed to possess a central, metastable bcc core [227]. In Penning traps, the stable fcc lattices observed in ion crystals with greater than 50000 ions [308, 155], which was discussed in the previous section, is also an example of a metastable configuration.
@@
[…]
II.3.3 Normal modes of three dimensional crystals
Normal mode phonon spectra of bulk, crystalline strongly coupled OCPs have been calculated with varying degrees of accuracy over the past 6 decades both in the absence  [60, 51, 9] and presence of a strong magnetic field [8, 322, 231]. Here, the interest has frequently been in predicting the thermodynamic properties of OCP such as dense astrophysical objects. As discussed in previous sections, trapped-ion crystals formed and studied in the laboratory typically are not of sufficient size to form a body-centered cubic lattice in a large fraction of the interior of the crystal. In addition, the normal modes that are most straight forwardly excited in the laboratory typically have wavelengths that are long compared to the inter-ion spacing. The dispersion relation of long wavelength modes is insensitive to the details of the lattice structure but can be influenced by the boundary or shape of the ion crystal.
A trap geometry that is frequently employed in experimental work provides quadratic (frequently called harmonic) confinement in the axial and radial directions. See Eqs. (13) and (22). As discussed in Secs. I.2.1 and I.2.2, at low laser cooling temperatures a quadratic confining potential results in uniform density crystals with a spheroidal shape. A spheroid is the shape obtained by rotating an ellipse about one of its principal axes. Interestingly, a simple exact analytic solution exists for the normal modes of a uniformly charged spheroid, both without and in the presence of a magnetic field. The treatment assumes that the normal mode wavelength is long compared to the inter-ion spacing so that the mode is well described by cold fluid equations. See Refs. [93, 28] for a detailed description of the cold-fluid equations.
The modes can be catalogued by two integers $l$ and $m$ where, following the conventions of 94, $0\leq m<l$. Here, $l$ and $m$ characterize the functional dependence of the mode potential $\psi$ in terms of associated Legendre functions. For example, the solution for the mode potential outside the spheroidal crystal is given by

### v2

@@ II.3.1 Ground state configuration
More surprisingly, when two different ion species are simultaneously confined in a linear rf quadrupole trap, cylindrical crystal structures can be realized that mimick those of infinitely long systems [159]. This should also be observable in Penning traps [204, 167]. See the paragraph on Two-species Coulomb crystals which concludes this subsection.
@@
Slab geometry— The slab geometry consists of a uniform density OCP that is harmonically confined in one direction, $q\phi_{T}(\rho,z)=m\omega_{z}^{2}z^{2}/2$, but infinite in the orthogonal directions. This geometry allows for a careful theoretical treatment that illustrates the importance of surface effects with 3D trapped ion crystals [97]. The treatment assumes that the lattice planes that form in the $T=0$ limit have the same symmetry. It is found that, although bcc has a lower bulk energy per ion than fcc, the fcc lattice has a smaller surface energy contribution than bcc, making it the lower energy lattice for thicknesses with smaller numbers of planes. Dubin showed that when the slab thickness is larger than approximately 60 planes, bcc becomes the structure that minimizes the energy, while for smaller thickness fcc prevails.
@@
Experimentally this geometry can be approximated in ion traps by adjusting the trap radial confinement to be much weaker than the axial confinement. In a Penning trap this condition can be achieved with low rotation frequencies $\omega_{r}$ that are slightly higher than the magnetron frequency $\omega_{m}$ of the trap (see Eq. (24) and Fig. 7). For a given rotation frequency, this is the regime that gives rise to single-plane crystals (see Sec. II.2) for a sufficiently small number of ions. With larger numbers of trapped ions, multi-plane crystals form with an overall pancake (or lenticular) shape. The lattice planes have some curvature, but near the radial center ($\rho\approx 0$) of the trap, the planes are very flat and can be compared with the structures of infinite planar systems analyzed by [100].
@@
Figure 32: Axial positions of the planes in the central region ($\rho\approx 0$) of small aspect ratio ion crystals as a function of the central areal number density. The lines show predictions from theory, and the symbols show experimental measurements. An open square (filled circle) indicates the experimental observation of a square or rhombic plane (hexagonal plane) was experimentally observed. Similarly, a thin line (thick line) indicates that theory predicts a square or rhombic plane (hexagonal plane). Lengths have been normalized by $a_{WS2D}=(3e^{2}/4\pi\epsilon_{0}m\omega_{z}^{2})^{1/3}$. From 238, where the interlattice displacement vector $\bf{c}_{2}$ is defined.
@@
Figure 32 summarizes the ion crystal structures observed in the laboratory as a function of the 2D areal number density ($\sigma_{0}a_{ws2D}^{2}$) obtained by integrating over the axial extent of the crystal [238]. The crystal structure was determined from direct images of the ion fluorescence from the central ($\rho\approx 0$) region of Doppler laser-cooled ion crystals with aspect ratios $\alpha\equiv Z_{p}/R_{p}<0.1$ and ion number $N\lesssim 10^{4}$ ions. Figure 32 also compares the measured separation of the planes with the theory of [100]. Starting with low 2D number density, the experiment reported single-plane crystals with a triangular lattice (see Sec.  II.2.2). With increasing $\sigma_{0}a_{ws2D}^{2}$, theory predicts a “buckling-transition” to three closely spaced planes that is stable over a very small interval of $\sigma_{0}a_{ws2D}^{2}$ (see Sec. II.2.4). This small region of three planes is followed by two, more widely separated planes. Initially each plane has a square and then rhombic symmetry. As the planes move further apart, they transition into planes with hexagonal symmetry. With increasing $\sigma_{0}a_{ws2D}^{2}$, eventually 3 planes with square symmetry form, and the pattern repeats. As visible from the figure, the overall agreement between experimental observations with the $T=0$, infinite extent theory of [100] is good, with some minor discrepancies [238]. One disagreement is that the transition from a single plane to a small region of closely spaced three planes was not observed. This transition is particularly interesting (see Sec. II.2.4) and motivates a more careful study at lower temperatures with new experimental tools.
@@
Interestingly, the crystal planes with rhombic symmetry, observed in [238], have peculiar thermodynamic properties such as negative Poisson ratios. A negative Poisson’s ratio means that stretching the crystal in a particular direction, in this case, along the $\hat{z}$ or axial direction of the trap, causes an expansion in a lateral direction. This behavior contradicts common experience for ordinary materials that act as incompressible, such as rubber bands. Negative Poisson ratios were measured in  [17].
@@
We close this section on slab geometries by pointing out that multilayer 3D crystals of trapped ions consisting of flat planes can be formed through the addition of anharmonic terms to the trapping potential. This was numerically demonstrated for crystals in Penning traps, where the anharmonic terms are additions to the usual harmonic trapping potential (see Eq. 20) and are readily implementable in the lab [150]. The formation and control of ion crystals in traps with significant anharmonic components to the trapping potential is a subject of ongoing experimental investigation.
@@
Two-species Coulomb crystals— As indicated previously, two-species Coulomb crystals constitute a very interesting and special type of Coulomb crystal. It is instructive to draw a comparison with normal solids formed through electronic (chemical) bindings between the atomic constituencies. In this case, when more atomic species are involved, various alloys form. In contrast, in Coulomb crystals the confined ions are mainly interacting through the repulsive Coulomb forces at a distance that inhibits chemical bonding. Consequently, the structural organization of two simultaneously trapped ion species is the one that minimizes the potential energy of all ions. Since in both linear Paul (rf) traps and Penning traps, the effective confining potential in the plane perpendicular to the static confining direction depends on both the charge and mass of a specific ion species, a two-species Coulomb crystal will generally lead to a separation of the two species with respect to this plane. This type of separation consists of the strongest bound species forming a cylinder-like central structure within an embedding spheroidal, globally shaped crystal and has been reported in [159, 254]. Examples are shown in Fig. 33 and Fig. 34. When the charge-to-mass ratio of the two species is the same, the two species mix. This has been observed in experiments where the two trapped species were 26Mg+ and 24MgD+ ions [242]. It was also reported in molecular dynamics simulations where one species has both double the mass and charge of the other [225]. In the latter case, the ground state of an infinite system is expected to be a simple cubic lattice with the unit cell consisting of one species at a corner and the other displaced half way the $\langle 111\rangle$ diagonal, similar to chemically bound CsCl crystals in solid state.
@@
Figure 33: (Color online) (a) Sketch of a linear rf trap experiment to study two-species ion Coulomb crystals consisting of Mg+ and Ca+ ions (a). The lower plots show fluorescence images for (b) 300 Mg+ ions (red central region) and 3000 Ca+ ions and (c) 47 Mg+ ions and 1300 Ca+ ions. Both crystals have a spheroidal global shape with a cylindrical symmetry axis coinciding with the central axis of the Mg+ component. From 159.
@@ II.3.2 Metastable and perturbation induced structures
The theoretical analysis of the experimental results discussed so far have assumed that the trapping potential is conservative and purely harmonic. However, this is only an approximation of the experimental situations. Specifically, in rf traps the time dependence of the trapping potential can produce measureable deviations from this idealized situation, thus impact the formation of ion crystal structures and even the resulting structures.
@@
For instance, in linear rf traps operated with $a_{\mathrm{dc}}=0$ the pseudopotential is symmetric under rotation about the $z$-axis in the geometry of Fig. 4. However, this continuous symmetry is only an approximation of the exact discrete symmetry, imposed by the four electrode configuration of the trap. The effect of this discrete symmetry is already evident in clusters composed of two ions, which are forced by a strong confining force to lie in the xyx-y plane: molecular dynamics simulations (as well as expansions of the pseudo potential beyond the leading order [92]) show that at equilibrium the two ions always align along one of the (1,1) or (1,-1) directions in the plane, where the orientation of the micromotion is perpendicular to the line connecting the two ions. Similarly, for weak confinement along the $z$-axis, small zigzag structures always align in the same directions.
@@
In three-dimensional ion crystals the micromotion occurs along several directions and the resulting order is therefore more challenging to predict. There are however some general properties extracted from experimental evidence. Three-dimensional crystals that are laser-cooled by beams propagating along the axis of the linear trap are never observed to rotate around the same axis. This occurs for single as well as for multi-species crystals and indicates that the trap potential tends to pin the crystals. Moreover, translationally periodic crystal structures of single species crystals seem to have a propensity to occur with, e.g., bcc or fcc configurations at specific orientations with respect to the trap electrodes [253, 254]. More pronounced effects have been observed in two-species crystals consisting of two isotopes of Ca+ (40Ca+ and 44Ca+), as shown in Fig. 34 [254]. Here, the inner 40Ca+ component, in addition to having a global cylindrical shape, had a very persistent internal structure (presumably orthorhombic) with fixed relation to the trap electrodes and with life times longer than 10 seconds. Similar results were found by simulations, which included the full rf trap potential [254].
@@
Figure 34: (Color online) Fluorescence images of two-species ion Coulomb crystals containing $\sim$1500 40Ca+ ions in the central region (red) and $\sim$2000 44Ca+ ions in the surrounding regions (blue) for the same ion densities, but different aspect ratios of the trapping potential. From 254.
The application of optically-induced dipole forces  [139] can generate spatially localized perturbations to the ions’ trapping potential. A standing-wave light field can force the ions to localize on gratings of potential minima separated by half the laser wavelength, $\lambda/2$ [67, 111, 216, 205, 370]. In general, the competition with the Coulomb repulsion gives rise to frustration (see Sec. III.5). However, for crystalline structures with bcc, fcc, or hcp symmetry, one could tune these gratings so that the periodicity is commensurate with the distance between the lattice planes. Molecular dynamics simulations showed that in this way one can control the crystal structures and orientations as well as adiabatically change the crystal structure from bcc to fcc structures, and vice versa [157]. In a cylindrically symmetric pseudopotential, the radiation pressure of a single traveling laser beam that does not propagate across the trap center exerts a torque and can set the crystal into rotation. In [93] this was observed by illuminating a crystal in a rf-trap by a beam propagating within the xyx-y plane, but off-set with respect to the center point. The measured rotational frequencies varied from zero all the way up to 98% of the radial trap frequency dependent on the specific position and detuning of the laser beam [93] (see Fig. 35 as well as Sec. III.5 for an analogous experiment performed in Penning traps).
Figure 35: Images of an ion crystal consisting of $\sim$200 24Mg+ ions illuminated by two traveling laser beams. One laser beam propagates along the the $z$-axis and laser cool the crystal. The other beam propagates within the xyx-y plane and exerts a torque on the crystal forcing a rotation, whose axis is left-right in the figure. The image in (a) corresponds to a stationary crystal. The other images show rotating crystals with increasing rotational frequency $\Omega_{\rm rot}$, namely: (b) $\Omega_{\rm rot}=0.4\omega_{\rm r,trap}$, (c) $\Omega_{\rm rot}=0.7\omega_{\rm r,trap}$, (d) $\Omega_{\rm rot}=0.9\omega_{\rm r,trap}$, (e) $\Omega_{\rm rot}=0.96\omega_{\rm r,trap}$ and (f) $\Omega_{\rm rot}=0.98\omega_{\rm r,trap}$, where $\omega_{\rm r,trap}=\omega_{y}=\omega_{x}$ is the radial trap frequency. In all experiments, $\omega_{\rm r,trap}=2\pi\times$180 kHz and $\omega_{z}=2\pi\times$86 kHz. From 93.
@@
In rf traps, a more dramatic kind of perturbation can be induced when the density of a crystal is increased to the point, where the plasma frequency of the crystal $\omega_{p}$ becomes close to the half that of the rf drive frequency, $\Omega_{\rm rf}$, namely, when $\omega_{p}\sim\Omega_{\rm rf}/2$. In such situations, the parametric resonances lead to heating of the crystal [29], and in many instances even to melting. However, if the crystal is set into rotation, its density and hence its plasma frequency will be lowered, and the crystal will eventually get out of parametric resonance with the rf frequency. By appropriate laser cooling and laser torque conditions, one can stabilize structures where the ions form rotating discs [191], see Fig. 36. A full understanding of these types of dynamics is still not available. Nevertheless these configurations seem to satisfy a detailed balance condition between the laser-cooling and the rf heating processes.
@@
As mentioned in Sec. II.3.1, the number of ordered metastable configurations within a given energy interval typically increases exponentially with the number of ions in a Coulomb crystal. Hence, in experiments where the ions are cooled to finite temperatures, one will expect to observe thermally excited metastable configurations. Figure 37 shows four images of the same cold ion ensemble of approximately 2700 ions stored in a linear rf trap. The images were recorded seconds apart, showing various 3D periodic lattices structures. These are all excited metastable configurations, e.g., they correspond to energies higher than the ground state energy. We note that Coulomb crystals as small as $\sim$1000 ions have been observed to possess a central, metastable bcc ordered core [253]. In Penning traps, examples of metastable structures are the fcc lattices observed in ion crystals with more than 50000 ions [343, 168], see the previous section.
@@
II.3.3 Normal modes of three-dimensional crystals
Normal mode phonon spectra of bulk, crystalline strongly coupled OCPs have been calculated with varying degrees of accuracy over the past 6 decades both in the absence  [65, 55, 10] and presence of a strong magnetic field [357, 257, 9]. Here, the interest has frequently been in predicting the thermodynamic properties of dense astrophysical OCPs. As discussed in previous sections, trapped-ion crystals formed and studied in the laboratory typically are not of sufficient size to form a body-centered cubic lattice in a large fraction of the interior of the crystal. In addition, the normal modes that are most straightforwardly excited in the laboratory typically have wavelengths that are long compared to the inter-ion spacing. The dispersion relation of long wavelength modes is insensitive to the details of the lattice structure but can be influenced by the boundary or shape of the ion crystal.
A trap geometry that is frequently employed in experimental work provides quadratic (frequently called harmonic) confinement in the axial and radial directions. See Eqs. (13) and (22). As discussed in Secs. I.2.1 and I.2.2, at low laser cooling temperatures a quadratic confining potential results in uniform density crystals with a spheroidal shape. Interestingly, a simple exact analytic solution exists for the normal modes of a uniformly charged spheroid, both in the absence as well as in the presence of a magnetic field. The treatment assumes that the normal mode wavelength is long compared to the inter-ion spacing so that the mode is well described by cold fluid equations. See Refs. [99, 31] for a detailed description of the cold-fluid equations.
The modes can be catalogued by two integers $l$ and $m$ where $0\leq m<l$, following the conventions of 100. Here, $l$ and $m$ characterize the functional dependence of the mode potential $\psi$ in terms of associated Legendre functions. For example, the solution for the mode potential outside the spheroidal crystal is given by